The cholesterol levels of women aged 21-40 in Canada are approximately Normally distributed with a mean of 190 miligrams per decilitre (mg/dl). In July of 2007, a clinical assessment applied in Toronto to random sample of twenty-nine Asian female immigrants aged 21-40 had a mean level of 179.52 mg/dl and a standard deviation of 38 mg/dl. (a) At the 10% level of significance test whether the mean cholesterol level of Asian women is the same as the national average. of Asian women is lower than the national average? all Canadian women aged 21-40 is 32.4 mg/dl. (b) With minor changes to the previous part, is there evidence that the mean cholesterol level (c) Repeat part (a) utilizing the fact that the standard deviation of the cholesterol levels for
Answer #1

a) 0 :i-190

H_{0}: \mu \neq 190

t=(\frac{179.52-190}{38})*\sqrt{29}

t = -1.485

P value = P(t<-1.485+P(t>1.485)

= 2*(1-T.DIST(1.485,28,TRUE))

= 0.1487

Since P value >0.10, we failed to conclude that mean cholestral level of asian women is different from national average

b) No, Since the 90% Interval will contain 0

c)

H_{0}: \mu =190

H_{0}: \mu \neq 190

z=(\frac{179.52-190}{32.4})*\sqrt{29}

z = -1.742

P value = P(z<-1.72+P(z>1.742)

2*NORMSDIST(G157)

= 0.0815

Since P value <0.10, we conclude that mean cholestral level of asian women is different from national average